OpenStudy (anonymous.):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. y = 3x - 12 y = -3x y = -3x + 12 y = 3x

OpenStudy (anonymous):

Do you know how to find the slope of a line?

OpenStudy (anonymous.):

Yea ! y-y^1=m(x-x^2) y-6=-3(x-2) ?

OpenStudy (anonymous):

Yep. So you know you can just look at that original equation and see the slope, right? slope-intercept form is y=mx+b. That makes the slope in your original equation... -3, right?

OpenStudy (anonymous.):

So it will be y = -3x + 12 ? :D

OpenStudy (anonymous):

You got it. :)

OpenStudy (anonymous.):

Thank you sooo much for your help :D

OpenStudy (anonymous):

No problem. :)